Find all exact solutions on the interval
0 ≤ x < 2𝜋. (Enter your answers as a comma-separated list.)
cot(x) + 7 = 6

Answer by math_tutor2020(3817)


cot(x) + 7 = 6
cot(x) = 6-7
cot(x) = -1
1/tan(x) = -1
tan(x) = -1

Use the unit circle to determine the solutions to tan(theta) = -1 are theta = 3𝜋/4 radians and theta = 7𝜋/4 radians.
These angles are in quadrants 2 and 4 respectively.

3𝜋/4 radians = 135 degrees
7𝜋/4 radians = 315 degrees
These two angles are separated by 180 degrees (aka 𝜋 radians), which is half a revolution.

Use a calculator in radian mode to confirm that
tan(3𝜋/4) = -1
tan(7𝜋/4) = -1
I'll let the student check each solution back into the original equation.


Answers: 3𝜋/4, 7𝜋/4